The length of a rectangle is 4cm more than twice its width. The perimeter is 86cm. Find the width and length.
Answer by tommyt3rd(5050)

1) l=4+2w
2) 2(l+w)=86
2) l+w=43
replacing l...
(4+2w)+w=43
3w=39
w=13
1) becomes
l=4+2*13
l=30